# Reading Recognition in the Wild

### v1

@@ 2 Related Work
(train/val/test) 81 people Libraries Patios Digital Short texts While walking Skimming Hard negatives sequences
Focus: diversity 1061 videos Homes Roads/trails Objects Non-texts Writing/typing Scanning (71%/29%) (reading /
Stores In the woods Dynamic texts Out loud not reading)
@@ 3.1 Overview
The dataset contains about $100$ hours of recordings of reading and non-reading activities collected from one RGB ($30$Hz, $1408$p, $110$° FoV) and two SLAM ($150$° FoV) cameras, two eye tracking cameras ($60$Hz, calibrated), two IMUs (with odometry outputs from visual SLAM), and audio transcribed using WhisperX [2]. We independently collect two subsets of this dataset, as in Table 1.
@@ 3.4 Data collection process
Privacy. We strictly followed Project Aria Research guidelines. All data has been de-identified, and faces and license plates were anonymized with EgoBlur [38].
@@ 4.4 Generalization
While we train on English texts, we find that our model generalizes well to other left-to-right languages across different writing systems, but struggles with vertical and right-to-left texts, as the gaze pattern is in a different direction. To address this, we find that simply augmenting the gaze at inference time (90° rotation for vertical texts and horizontal flip for right-to-left texts) allows the model to generalize well. In practical scenarios, this can be done depending on geo-location.
@@ 6.3 Generalization
Cross-dataset generalization. To demonstrate the importance of collecting reading examples in freeform settings, we conduct experiments to test for generalizability across datasets. For this, we utilize EGTEA Gaze+ dataset [29], where we only use their ‘reading’ action labels, and treat other labels as not reading. To match the data available in EGTEA, we use 2D gaze projection at $30$Hz. We conduct cross-generalization experiments where we train on one training set and evaluate on the other test set. We show that training on EGTEA with limited training samples does not generalize to in-the-wild scenarios, whereas the generalization gap for our dataset is much smaller.
Gaze RGB IMU Acc F1 Latency
@@
1 No read 0.88 0.04 0.02 0.02 0.01 0.03 0.00
2 Walk 0.09 0.85 0.04 0.01 0.00 0.00 0.01
3 Out loud 0.13 0.02 0.64 0.17 0.02 0.01 0.01
4 Engaged 0.14 0.02 0.06 0.54 0.12 0.01 0.11
5 Scan 0.08 0.01 0.03 0.39 0.41 0.00 0.08
6 Write/type 0.49 0.01 0.03 0.02 0.05 0.39 0.01
7 Skim 0.13 0.04 0.05 0.47 0.15 0.00 0.16
@@ 6.5 Ablation studies
Model size. We experimented other model sizes, with XS, S, M, L having 8, 16, 32, and 64 latent dimensions respectively. We also experimented with a pretrained image encoder (MobileNetV3-S) in the XL variant. We find that stronger model results in better results, and notably the S model performs surprisingly well with only $6$k parameters.

### v3

@@ 2 Related Work
(train/val/test) 81 people Libraries Patios Digital Short texts Walking Skimming Hard negatives sequences
Focus: diversity 1061 videos Homes Roads/trails Objects Non-texts Writing Scanning (71%/29%) (reading /
Stores In the woods Dynamic texts Typing Out loud not reading)
@@ 3.1 Overview
The dataset contains about $100$ hours of recordings of reading and non-reading activities collected from one RGB (3030Hz, 14081408p, 110110° FoV) and two SLAM (150150° FoV) cameras, two eye tracking cameras (6060Hz, calibrated), two IMUs (with odometry outputs from visual SLAM), and audio transcribed using WhisperX [2]. We independently collect two subsets of this dataset, as in Table 1.
@@ 3.4 Data collection process
Privacy. We strictly followed Project Aria Research guidelines. All data has been de-identified, and faces and license plates were anonymized with EgoBlur [35]. We source the venues ourselves do not use the participants’ private spaces to prevent exposure of sensitive or identifiable information.
@@ 4.4 Generalization
While we train on English texts, we find that our model generalizes well to other left-to-right languages across different writing systems, but struggles with vertical and right-to-left texts, as the gaze pattern is in a different direction. To address this, we find that simply augmenting the gaze at inference time (90° rotation for vertical texts and horizontal flip for right-to-left texts) allows the model to generalize well. In practical scenarios, this can be done depending on geo-location. During training, we also add a small fraction of rotated gaze to help with reading vertical texts.
@@ 6.3 Generalization
Further, we also notice the differences in reading speed across different users, especially across different languages. It is possible to personalize the model by scaling the gaze to the magnitude of the reader, and empirically this solves some of the failure cases.
@@
Cross-dataset generalization. To demonstrate the importance of collecting reading examples in freeform settings, we conduct experiments to test for generalizability across datasets. For this, we utilize EGTEA Gaze+ dataset [27], where we only use their ‘reading’ action labels, and treat other labels as not reading. To match the data available in EGTEA, we use 2D gaze projection at 3030Hz. We conduct cross-generalization experiments where we train on one training set and evaluate on the other test set. We show that training on EGTEA with limited training samples does not generalize to in-the-wild scenarios, whereas the generalization gap for our dataset is much smaller.
Gaze RGB IMU Acc F1 Latency (s)
@@
[…]
@@ 6.4 Application: real-time reading detection
Practical deployment. We also investigate whether model of such size can be run practically. From parallel comparisons, the model can indeed comfortably run real-time on Aria Gen 2 glasses on-device, without the need to off-load the model to online computation. Given the estimated power consumption, the glasses can run for at least 4 hours continuously (inclusive of the base power consumption for basic computation, power delivery and sensor suites, and the thermal constraints). As the model runs on-device without having to send the model input and output back and forth to the server (as would have been done with, say, VLMs), the latency is negligible.
@@ 6.5 Ablation studies
Model size. We experimented other model sizes, with XS, S, M, L having 8, 16, 32, and 64 latent dimensions respectively. We also experimented with a pretrained image encoder (MobileNetV3-S) in the XL variant. We find that stronger model results in better results, and notably the S model performs surprisingly well with only 66k parameters.
@@ 6.6 Extension: understanding types of reading
7 Broader Impact
Always-on smart glasses raise important questions about social acceptability, both for the wearer and for the public, especially when such technologies are deployed at scale. We hope that the ideas presented in this paper can also help mitigate such concerns.
Safety. Sensitive personal data, such as eye gaze, introduces unique risks. Our algorithm runs fully on-device, which ensures that sensitive information does not need to leave the user’s device. This is a step toward stronger privacy protections for wearers. At the same time, we acknowledge that using eye gaze as a signal creates new challenges. Eye movement can reveal intentions, interests, and even emotional states, which raises a distinct category of privacy concerns.
Surveillance. Our work aims to reduce reliance on invasive sensing. First, by leveraging eye gaze, we minimize the required front-camera capture to a very small patch (0.2% of the full image) rather than recording the entire scene. Second, our approach can operate solely on eye gaze data without requiring any camera input. More broadly, eye gaze offers a powerful cue about where the user is looking, which enables RGB capture to be more targeted. This reduces the risk of collecting unintended or intrusive information about bystanders. We hope that future algorithms continue in this direction.
Data Governance. We follow the Project Aria Research Guidelines and will release our system with a Responsible Use Policy to promote ethical research practices and to support safe deployment.